For the system shown in Fig. 2, X(e), the Fourier transform of the input signal x[n], is shown in Fig. 3 5.

Interpolator Decimator Lowpass filter Gain = L Lowpass filter In Sampling - period: T TM Lowpass filter Gain = L [n を[n] | . Cutoff= min(mIL, TIM) aln Sampling period: TM Fig. 2. For Problems 4 and 5

Fig. 3. For Problem 5 For each of the following choices of L and M, specify the maximum possible value of oo such that Xa (ee) = ax (eMel) for some constant a. (a) M 5, L 3 (b) M-3, L- 5
Answer #1

Given:

Here in the problem the given signal x[n] in first interpolated by 'L' followed by decimation by 'M' to obtain xd[n], with overall cut-off frequency (let wc) :

From the scaling property of DTFT:

If,

then, , for L> 1 i.e Interpolation

And, if x[n] is decimated by M, then,

Thus,overall,

, where, k= 0,1,2,3......(eqn.A)

Now let us take the problem for different values of L and M given in the question.

a) M =5 and L=3

Since, M > L, the cut-off frequency will be:

for the signal xd[n]:

Thus, from eqn A, the first pulse of xd [n] lies between (-3w/5, 3w/5) for K=0,

This implies,  ,

Thus,

.

b) M=3 and L=5.

Since, L > M, the cut-off frequency will be:

for the signal xd[n]:

Thus, from eqn A, the first pulse of xd [n] lies between (-5w/3, 5w/3) for K=0,

This implies, ,

Thus,

.